Joe and Mary working together can rake a lawn in 2 hours. Joe can do the job alone in 3 hours. How long would it take Mary to rake the lawn alone?
Answer by stanbon(48579)

Joe and Mary working together can rake a lawn in 2 hours. Joe can do the job alone in 3 hours. How long would it take Mary to rake the lawn alone?
Joe Data:
time to do job=3 hrs.; rate = 1/3 job per hr.
Mary Data:
time to do job=x hrs.; rate = 1/x job per hr.
Together Data:
time to do job=2 hrs; rate = 1/2 job per hr.
EQUATION:
Joe rate + Mary rate = Together rate
1/3 + 1/x = 1/2
Multiply thru by 6x to get:
2x + 6 = 3x
x=6 hrs. (time is would take Mary alone)
Cheers,
Stan H.